OpenStudy (anonymous):

Solve the polynomial equation by factoring and then using the zero product principle. 3x4 - 300x2 = 0

OpenStudy (anonymous):

A) {-10, 0, 10} B) {-10, 10} C) {0} D) \[-10\sqrt{3},0,10\sqrt{3}\]

OpenStudy (anonymous):

First you must factor out what you can. So in this case you can factor out a 3x^2. What do you get when you do that?

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

oh thats suppose to be 3x^4

OpenStudy (anonymous):

Yeah so factor out 3x^2 from 3x^4-300x^2=0 You should get 3x^2(x^2-100)=0

OpenStudy (anonymous):

Now if you have a times b=0 or ab=0 then either a=0 or b=0 or they both = 0 Now you can set 3x^2=0 and (x^2-100)=0 and solve for x

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

Well if 3x^2=0, then x has to equal 0 for one of the solutions. Then for the second equation x^2-100=0 use algebra to solve for x. Add 100 to both sides and then take the square root of both sides. Remember when you take a square root you will get a positive and negative answer.

OpenStudy (anonymous):

so for x=-10

OpenStudy (anonymous):

and 3x^2=0 = 0

OpenStudy (anonymous):

x is going to equal 10 and -10 for x^2-100=0 and x is equal to 0 for 3x^2=0

OpenStudy (anonymous):

ok thank you the answer is a?

OpenStudy (anonymous):

yes